OpenStudy (anonymous):

Solve. 7 – 3x = 3(3 – x) – 2

OpenStudy (anonymous):

7-3x=3(3-x)-2 =>7-3x=9-3x-2 =>7-3x=7-3x So, for any value of x both side are equal

OpenStudy (anonymous):

so they are real numbers?

OpenStudy (anonymous):

yes they are real but only when value of x is real